# Online Learning with Probing for Sequential User-Centric Selection

### v1

@@ Abstract
We formalise sequential decision–making with information acquisition as the Probing-augmented User-Centric Selection (PUCS) framework, where a learner first probes a subset of arms to obtain side information on resources and rewards, and then assigns $K$ plays to $M$ arms. PUCS encompasses practical scenarios such as ridesharing, wireless scheduling, and content recommendation, in which both resources and payoffs are initially unknown and probing incurs cost. For the offline setting (known distributions) we present a greedy probing algorithm with a constant-factor approximation guarantee of $\zeta=(e-1)/(2e-1)$. For the online setting (unknown distributions) we introduce OLPA, a stochastic combinatorial bandit algorithm that achieves a regret bound of $\mathcal{O}\!\bigl{(}\sqrt{T}+\ln^{2}\!T\bigr{)}$. We also prove an $\Omega(\sqrt{T})$ lower bound, showing that the upper bound is tight up to logarithmic factors. Numerical results using the dataset collected from the real world demonstrate the effectiveness of our solutions.
@@ 3.1 Arm, Play, and Resource Model
Consider a sequential user-centric decision problem [7] over $T\in\mathbb{N}^{+}$ time slots, which is a variant of the classic multi-play multi-armed bandits problem. There are a set of $M\in\mathbb{N}^{+}$ arms and a set of $K\in\mathbb{N}^{+}$ plays. In each time slot, the decision is to assign the $K$ plays denoted by $[K]:=\{1,2,\ldots,K\}$ to the arm set denoted by $[M]:=\{1,2,\ldots,M\}$. Each play can only be assigned to one arm and different plays can be assigned to the same arm.
@@ 3.2 Probing Model
::=:= $\mathbb{E}\left[\sum_{i=1}^{j}{\bm{\mu}}_{m,i}^{\text{sort}}\right],~j:=\min\{D_{t,m},|C_{t,m}|\}$
@@ 3.3 Objective and Regret
:=:= $\max\limits_{\begin{subarray}{c}C_{m}\subseteq[K]\\ m\in[M]\end{subarray}}\left(\sum_{m\in M\setminus S}\mathcal{R}_{m}(C_{m};{\bm{\mu}}_{m},{\mathbf{p}}_{m})\right.$
@@ 4.1 Offline Greedy Probing
For small problems, the optimal solution for the offline setting can be obtained by solving Problem (2) with an exhaustive search. For larger instances, however, a more efficient solution is needed. In this section, we derive a greedy algorithm that provides a constant factor approximation. One work [24] states that a simple greedy algorithm with an objective function that is monotone and submodular can obtain an approximation factor of an approximation factor of $1-1/e$. Based on this, we construct our objective function $f_{\text{prob}}(S)$ that is monotone and submodular shown in Lemma 3 and Lemma 4 and design an Offline Greedy Probing algorithm to obtain the approximation solution.
@@
[…]
We note that if $S=\emptyset$, $R(S)=f(S)=f_{\text{unprobed}}(S)$. With these notations, we are ready to explain our greedy probing algorithm for the offline setting (Algorithm 1). It starts by initializing $I$ empty sets $S_{i}$ for $i=0$ to $I-1$, where $I$ is the probing budget (lines 1-2). In each iteration from 1 to $I-1$, the algorithm selects the arm $m$ that, when added to the current set $S_{i-1}$, provides the maximum marginal increase in the expected reward function $f_{\text{prob}}$ (line 4). The algorithm then updates the probing set $S_{i}$ to include the newly selected arm (line 5). It determines the optimal probing set $S^{\rm pr}$ by selecting the set that maximizes the adjusted reward function $(1-\alpha(i))f_{\text{prob}}(S_{i})$, where $\alpha(i)$ captures the probing cost as a function of the set size (lines 6-7). If the adjusted reward for the selected probing set $S_{j}$ is less than the reward obtained without probing, $f_{\text{unprobed}}(\emptyset)$, then the final probing set $S^{\rm pr}$ is set to be empty (lines 8-9).
@@ 5.1 Parameters
We use history results from time step 111 to $t$ to update our empirical parameters. Specifically, we define the empirical probability mass function $\hat{p}_{m,d}^{(t)}=\sum_{i=1}^{t}\textbf{1}\{D_{m,i}=d\}/t,\forall m\in[M],d\in[D_{\text{max}}]$ and $n^{(t)}_{m,k}:=\sum_{s=1}^{t}\mathbf{1}\{R_{i,m,k}\neq\text{null}\}\mathbf{1}\{k\in C_{i,m}\},\forall m\in[M],k\in[K]$ and empirical mean $\hat{\mu}^{(t)}_{m,k}\leftarrow\sum_{i=1}^{t}\mathbf{1}\{R_{i,m,k}\neq\text{null}\}R_{i,m,k}\mathbf{1}\{k\in C_{i,m}\}/n^{(t)}_{m,k},\forall m\in[M],k\in[K]$ and empirical CDF $\hat{F}^{(t)}_{m,k}({\mathbf{x}})\leftarrow\sum_{i=1}^{t}\mathbf{1}\{R_{i,m,k}\neq\text{null}\}\mathbf{1}\{R_{i,m,k}\leq x\}\mathbf{1}\{k\in C_{i,m}\}/n^{(t)}_{m,k},\forall x\in\mathbb{R},m\in[M],k\in[K]$
@@ 6 Experiments
The reward is based on the Manhattan distance between a vehicle and a pickup location. The distances are normalized to the [0, 1] range, where closer distances correspond to higher rewards. We consider two types of reward distributions: (1) a Bernoulli distribution with mean $\mu_{m,k}$, representing the normalized reward for arm $m$ and play $k$, assumed i.i.d. across $t$, $m$, and $k$; and (2) a discrete distribution with support [0.1,0.4,0.7,1.0][0.1,0.4,0.7,1.0][ 0.1 , 0.4 , 0.7 , 1.0 ], where probabilities for each level are derived from the distribution of normalized rewards. Cumulative regret is computed using Equation (3), where $R(S^{\star})$ is obtained through exhaustive search. Specific details are provided in the supplementary material.
@@ 6.2 Results
Table 1 reports cumulative regret at 1 k, 2 k, and 3 k steps on the Chicago Taxi Trips dataset. In every configuration OLPA attains the smallest regret—e.g. in setting (b) it reduces regret by over 30% compared to Non-Probing and by a factor of 3$+$ relative to GR/RR—confirming that probing yields large constant-level gains in practice.
@@ 7 Conclusion
In this section, we present the technical proofs that support the results of our paper and details of the experimental setup. The structure is as follows:
  • Section 1: We prove the total expected reward for arm $m$, as defined in the Probing-Augmented User-Centric Selection (PUCS) Framework of our main paper.
@@
  • Section 5: This section presents the details of the experimental setup and data processing.
Appendix 1 Proof of the Total Expected Reward for Arm $m$
In the problem formulation, the total expected reward returned by arm $m$ is given by:
$\mathcal{R}_{m}(C_{t,m};{\bm{\mu}}_{m},{\mathbf{p}}_{m})$ $:=\mathbb{E}\left[\sum_{i=1}^{j}{\bm{\mu}}_{m,i}^{\text{sort}}\right],\quad\text{where }j=\min\{D_{t,m},|C_{t,m}|\}$
Proof.
The total expected reward $\mathcal{R}_{m}(C_{t,m};{\bm{\mu}}_{m},{\mathbf{p}}_{m})$ can be understood as follows:
1. Case 1: Number of Resources $\geq$ Number of play
If the number of resources $D_{t,m}$ is greater than or equal to the number of plays $|C_{t,m}|$ assigned to arm $m$, the reward is simply the sum of the rewards received by these plays. In this case, every play in $C_{t,m}$ is matched with a resource, and the total reward is the sum of all these rewards.
2. Case 2: Number of Resources $<$ Number of play
If $D_{t,m}<|C_{t,m}|$, not all plays in $C_{t,m}$ can receive resources. The optimal approach is to allocate the limited resources to the plays that have the highest rewards. Thus, we sort the plays by their rewards in descending order and match resources to the top $D_{t,m}$ plays.
The expected reward, considering both cases, can be calculated as follows:
$\mathbb{E}\left[\sum_{i=1}^{j}{\bm{\mu}}_{t,m,i}^{\text{sort}}\right],\quad j:=\min\{D_{t,m},|C_{t,m}|\}$ $=\sum_{d=1}^{|C_{t,m}|}\left(p_{m,d}\sum_{i=1}^{d}{\bm{\mu}}_{t,m,i}^{\text{sort}}\right)$
$\quad+\sum_{d=|C_{t,m}|+1}^{D_{\max}}p_{m,d}\left(\sum_{i=1}^{|C_{t,m}|}{\bm{\mu}}_{t,m,i}^{\text{sort}}\right).$
Here’s why this formula works:
- First Term: For $d\leq|C_{t,m}|$, every possible resource allocation scenario is considered. The probability of having exactly $d$ resources available is $p_{m,d}$, and the total reward is the sum of the top $d$ rewards in $C_{t,m}$.
- Second Term: For $d>|C_{t,m}|$, all plays in $C_{t,m}$ receive resources. The reward in this scenario is the sum of all rewards in $C_{t,m}$, and the probability is still $p_{m,d}$.
We compute the expected reward by integrating over all possible values of $D_{t,m}$ (from 1 to $D_{\max}$):
$\mathbb{E}\left[\sum_{i=1}^{j}{\bm{\mu}}_{t,m,i}^{\text{sort}}\right]$ $=\sum_{d=1}^{|C_{t,m}|}p_{m,d}\sum_{i=1}^{d}{\bm{\mu}}_{t,m,i}^{\text{sort}}+\sum_{d=|C_{t,m}|+1}^{D_{\max}}p_{m,d}\sum_{i=1}^{|C_{t,m}|}{\bm{\mu}}_{t,m,i}^{\text{sort}}$
$=\sum_{d=1}^{|C_{t,m}|}\left(p_{m,d}\sum_{i=1}^{d}{\bm{\mu}}_{t,m,i}^{\text{sort}}\right)+\sum_{d=|C_{t,m}|+1}^{D_{\max}}p_{m,d}\left(\sum_{i=1}^{|C_{t,m}|}{\bm{\mu}}_{t,m,i}^{\text{sort}}\right).$
@@ Appendix 3 Proofs for the offline case
$C^{(1)}_{m}:=\begin{cases}C^{*,\text{total}}_{m},~\text{ if $m\in S$}\\ \emptyset,~\text{ if $m\notin S$}\end{cases},~~C^{(2)}_{m}:=\begin{cases}C^{*,\text{total}}_{m},~\text{ if $m\in[M]\setminus S$}\\ \emptyset,~\text{ if $m\in S$}\end{cases},$
@@
then ${\mathbf{C}}^{(3)}$ satisfying the constraint of optimization problems defining $f_{\text{unprobed}}(S,\{{\bm{\mu}}_{m},{\mathbf{p}}_{m}\}_{m\in[M]})$, since $[M]\setminus T\subseteq[M]\setminus S$. It follows that
@@
$C^{(4)}_{m}:=\begin{cases}C^{*,S}_{m},~\text{ if $m\in S$}\\ \emptyset,~\text{ if $m\notin S$}\end{cases},$
@@
To show $f_{\text{prob}}(S)$ is submodular [24], we need to show that for any $S,T\subseteq[M]$, $f_{\text{prob}}(S)+f_{\text{prob}}(T)\geq f_{\text{prob}}(S\cap T)+f_{\text{prob}}(S\cup T)$
@@
[…]
@@
$R(S^{\text{pr}})=f(\emptyset)=f_{\text{unprobed}}(\emptyset)>(1-\alpha(|\tilde{S}^{*}|))f_{\text{prob}}(\tilde{S}^{*}).$
@@
The first inequality follows by Lemma (1); the second inequality follows by $1-\alpha(|S^{*}|)\leq 1$; the third inequality follows by and Lemma 2; the fourth inequality follows by [24], which states that a simple greedy algorithm with an objective function that is monotone and submodular can obtain an approximation factor of an approximation factor of $1-1/e$. Since $f_{\text{prob}}(S)$ is monotonely increasing and submodular shown in Lemma 3 and Lemma 4 respectively, we have the approximation factor $1-1/e$. The next one is the second equality, and it follows by $|S_{i}|=i$ for all $0\leq i\leq I$; the fifth inequality by the fact that $\tilde{S}^{*}$ maximizes $(1-\alpha(|S_{i}|))f\left(S_{i}\right)$; the last inequality follows by Lemma 5. ∎
@@ Appendix 4 Proofs for the online case
$=$ $\left(1-\alpha(|S|)\right)\left|\left(\sum_{m\in S}\mathcal{R}_{m}^{\text{prob}}(C_{m};{\mathbf{X}}_{m},N_{m})+\sum_{m\in[M]\setminus S}\mathcal{R}_{m}(C_{m};{\bm{\mu}}_{m},{\mathbf{p}}_{m})\right)-\right.$
@@
$=$ $\left(1-\alpha(|S|)\right)\left[\left(\sum_{m\in S}\mathcal{R}_{m}^{\text{prob}}(C_{m};{\mathbf{X}}_{m},N_{m})+\sum_{m\in[M]\setminus S}\mathcal{R}_{m}(C_{m};\bar{{\bm{\mu}}}_{m},{\mathbf{p}}_{m})\right)-\right.$
@@
where $\delta\in(0,1)$, $\epsilon^{(t)}_{m,k}=\sqrt{\left(1+{n^{(t)}_{m,k}}\right)\frac{\ln\left(\sqrt{n^{(t)}_{m,k}+1}/\delta\right)}{2\left(n^{(t)}_{m,k}\right)^{2}}}$ if $n^{(t)}_{m,k}>0$ and $\epsilon^{(t)}_{m,k}=+\infty$ if $n^{(t)}_{m,k}=0$.
@@
[…]
@@
[…]
@@
[…]
@@
Appendix 5 Experimental Setup and Data Processing
Datasets
We utilized two real-world datasets:
  • NYYellowTaxi 2016 Dataset: Contains trip records, including pickup and dropoff locations, passenger counts, and trip durations in New York City from March 22, 2016, to March 31, 2016.
  • Chicago Taxi Trips 2016 Dataset: A subset of data spanning from January 9, 2016, to September 29, 2016, was randomly selected. These datasets enable evaluation across diverse urban environments.
Data Preprocessing
  1. 1.
    Discretization of Pickup Locations and Cell Mapping: To discretize the geographical region, the latitude and longitude ranges of the pickup locations were divided into uniform bins of size 0.010.010.01 degrees. The center of each bin represents the center of a grid cell. Each pickup location was assigned to its nearest grid cell based on the geographic proximity to the cell center.
  2. 2.
    Passenger Probability Distribution (PMF): Within each grid cell, the total number of passenger pickups was calculated. The frequency of passenger counts in each cell was normalized to derive the probability mass function (PMF) for each arm:
    • Each grid cell corresponds to an arm $m$.
    • Let $c_{m,d}$ represent the count of pickups in grid cell $m$ with $d$ passengers. The total pickup count across all cells is denoted as $C=\sum_{m}\sum_{d}c_{m,d}$.
    • The PMF for arm $m$ is defined as ${\mathbf{p}}_{m}=[p_{m,d}]_{d=1}^{D_{\max}}$, where:
      $p_{m,d}=\frac{c_{m,d}}{C}.$
    • Grid cells with zero pickups were excluded from the PMF to ensure validity.
    The PMF was used to model the resource $D_{t,m}$ for arm $m$, which was sampled i.i.d. over time.
  3. 3.
    Vehicle and Play Mapping: Each vehicle represents a play $k$, and its location was randomly sampled within the dataset’s bounds. Vehicle locations were fixed throughout the experiment. Each pickup location is mapped to an arm (grid cell), and rewards for a given vehicle-pickup pair are computed based on their respective grid cells and locations.
Reward Modeling
  1. 1.
    Reward Calculation Based on Distance: Manhattan distances between vehicles and pickup locations were computed as weighted sums of latitude and longitude differences, converted to kilometers. The weights were defined as 111 km per degree for latitude and 85 km per degree for longitude. Rewards were computed as the negative of normalized distances, scaled to the range [0,1][0,1][ 0 , 1 ], using the formula:
    $\text{reward}=1-\frac{d-d_{\min}}{d_{\max}-d_{\min}}$
    where $d_{\min}$ and $d_{\max}$ are the minimum and maximum distances across all vehicle-pickup pairs. Smaller distances corresponded to higher rewards.
  2. 2.
    Reward Types:
    • Bernoulli Distribution: Rewards followed a Bernoulli distribution with mean $\mu_{m,k}$, where $\mu_{m,k}$ is the reward value computed using the formula above for arm $m$ and play $k$. These rewards were assumed i.i.d. across time steps, arms, and plays.
    • General Discrete Distribution: Rewards were discretized into levels with support [0.1,0.4,0.7,1.0][0.1,0.4,0.7,1.0][ 0.1 , 0.4 , 0.7 , 1.0 ]. The probabilities for each reward level were derived from the distribution of normalized rewards across all vehicle-pickup pairs:
      • For each vehicle and pickup location pair, the normalized reward was calculated using the Manhattan distance and scaled to the range [0,1][0,1][ 0 , 1 ].
      • The range [0,1][0,1][ 0 , 1 ] was divided into equal intervals corresponding to the discrete reward levels (e.g., [0,0.25)[0,0.25)[ 0 , 0.25 ) maps to 0.10.10.1, [0.25,0.5)[0.25,0.5)[ 0.25 , 0.5 ) maps to 0.40.40.4, etc.).
      • For each interval, the proportion of normalized rewards falling within that interval was computed. These proportions served as the probabilities for the corresponding reward levels.

### v2

@@ Abstract
We formalize sequential decision–making with information acquisition as the Probing-augmented User-Centric Selection (PUCS) framework, where a learner first probes a subset of arms to obtain side information on resources and rewards, and then assigns $K$ plays to $M$ arms. PUCS encompasses practical scenarios such as ridesharing, wireless scheduling, and content recommendation, in which both resources and payoffs are initially unknown and probing incurs cost. For the offline setting (known distributions) we present a greedy probing algorithm with a constant-factor approximation guarantee of $\zeta=(e-1)/(2e-1)$. For the online setting (unknown distributions) we introduce OLPA, a stochastic combinatorial bandit algorithm that achieves a regret bound of $\mathcal{O}\!\bigl{(}\sqrt{T}+\ln^{2}\!T\bigr{)}$. We also prove an $\Omega(\sqrt{T})$ lower bound, showing that the upper bound is tight up to logarithmic factors. Numerical results using the dataset collected from the real world demonstrate the effectiveness of our solutions.
@@ 3.1 Arm, Play, and Resource Model
Consider a sequential user-centric decision problem [7] over $T\in\mathbb{N}^{+}$ time slots, which is a variant of the classic multi-play multi-armed bandits problem. There are a set of $M\in\mathbb{N}^{+}$ arms and a set of $K\in\mathbb{N}^{+}$ plays. In each time slot, the decision is to assign the $K$ plays denoted by $K]:=\{1,2,\ldots,K\}$ to the arm set denoted by $M]:=\{1,2,\ldots,M\}$. Each play can only be assigned to one arm and different plays can be assigned to the same arm.
@@ 3.2 Probing Model
$:=$ $\mathbb{E}\left[\sum_{i=1}^{j}{\bm{\mu}}_{m,i}^{\text{sort}}\right],~j:=\min\{D_{t,m},|C_{t,m}|\}$
@@ 3.3 Objective and Regret
= $\max\limits_{\begin{subarray}{c}C_{m}\subseteq[K]\\ m\in[M]\end{subarray}}\left(\sum_{m\in M\setminus S}\mathcal{R}_{m}(C_{m};{\bm{\mu}}_{m},{\mathbf{p}}_{m})\right$.
@@ 4.1 Offline Greedy Probing
For small problems, the optimal solution for the offline setting can be obtained by solving Problem (2) with an exhaustive search. For larger instances, however, a more efficient solution is needed. In this section, we derive a greedy algorithm that provides a constant factor approximation. One work [24] states that a simple greedy algorithm with an objective function that is monotone and submodular can obtain an approximation factor of an approximation factor of 11/e1-1/e. Based on this, we construct our objective function $f_{\text{prob}}(S)$ that is monotone and submodular shown in Lemma 3 and Lemma 4 and design an Offline Greedy Probing algorithm to obtain the approximation solution.
@@
$\text{ s.t. }~C_{m_{1}}\cap C_{m_{2}}=\emptyset,~\forall m_{1}\neq m_{2},m_{1},m_{2}\in[M]\setminus S$
We note that if $S=\emptyset$, $R(S)=f(S)=f_{\text{unprobed}}(S)$. With these notations, we are ready to explain our greedy probing algorithm for the offline setting (Algorithm 1). It starts by initializing $I$ empty sets $S_{i}$ for $i=0$ to I1I-1, where $I$ is the probing budget (lines 1-2). In each iteration from 1 to I1I-1, the algorithm selects the arm $m$ that, when added to the current set $S_{i-1}$, provides the maximum marginal increase in the expected reward function $f_{\text{prob}}$ (line 4). The algorithm then updates the probing set $S_{i}$ to include the newly selected arm (line 5). It determines the optimal probing set $S^{\rm pr}$ by selecting the set that maximizes the adjusted reward function $(1-\alpha(i))f_{\text{prob}}(S_{i})$, where $\alpha(i)$ captures the probing cost as a function of the set size (lines 6-7). If the adjusted reward for the selected probing set $S_{j}$ is less than the reward obtained without probing, $f_{\text{unprobed}}(\emptyset)$, then the final probing set $S^{\rm pr}$ is set to be empty (lines 8-9).
@@ 5.1 Parameters
We use history results from time step $1$ to $t$ to update our empirical parameters. Specifically, we define the empirical probability mass function $\hat{p}_{m,d}^{(t)}=\sum_{i=1}^{t}\textbf{1}\{D_{m,i}=d\}/t,\forall m\in[M],d\in[D_{\text{max}}]$ and $n^{(t)}_{m,k}:=\sum_{s=1}^{t}\mathbf{1}\{R_{i,m,k}\neq\text{null}\}\mathbf{1}\{k\in C_{i,m}\},\forall m\in[M],k\in[K]$ and empirical mean $\hat{\mu}^{(t)}_{m,k}\leftarrow\sum_{i=1}^{t}\mathbf{1}\{R_{i,m,k}\neq\text{null}\}R_{i,m,k}\mathbf{1}\{k\in C_{i,m}\}/n^{(t)}_{m,k},\forall m\in[M],k\in[K]$ and empirical CDF $\hat{F}^{(t)}_{m,k}({\mathbf{x}})\leftarrow\sum_{i=1}^{t}\mathbf{1}\{R_{i,m,k}\neq\text{null}\}\mathbf{1}\{R_{i,m,k}\leq x\}\mathbf{1}\{k\in C_{i,m}\}/n^{(t)}_{m,k},\forall x\in\mathbb{R},m\in[M],k\in[K]$
@@ 6 Experiments
The reward is based on the Manhattan distance between a vehicle and a pickup location. The distances are normalized to the [0, 1] range, where closer distances correspond to higher rewards. We consider two types of reward distributions: (1) a Bernoulli distribution with mean $\mu_{m,k}$, representing the normalized reward for arm $m$ and play $k$, assumed i.i.d. across $t$, $m$, and $k$; and (2) a discrete distribution with support $[0.1,0.4,0.7,1.0]$, where probabilities for each level are derived from the distribution of normalized rewards. Cumulative regret is computed using Equation (3), where $R(S^{\star})$ is obtained through exhaustive search. Specific details are provided in the supplementary material.
@@ 6.2 Results
Table 1 reports cumulative regret at 1 k, 2 k, and 3 k steps on the Chicago Taxi Trips dataset. In every configuration OLPA attains the smallest regret—e.g. in setting (b) it reduces regret by over 30% compared to Non-Probing and by a factor of 3++ relative to GR/RR—confirming that probing yields large constant-level gains in practice.
@@ 7 Conclusion
In this section, we present the technical proofs that support the results of our paper. The structure is as follows:
@@ Appendix 2 Proofs for the offline case
[…]
@@
then ${\mathbf{C}}^{(3)}$ satisfying the constraint of optimization problems defining $f_{\text{unprobed}}(S,\{{\bm{\mu}}_{m},{\mathbf{p}}_{m}\}_{m\in[M]})$, since $M]\setminus T\subseteq[M]\setminus S$. It follows that
@@
[…]
@@
To show $f_{\text{prob}}(S)$ is submodular [24], we need to show that for any $S,T\subseteq[M]$, $f_{\text{prob}}(S)+f_{\text{prob}}(T)\geq f_{\text{prob}}(S\cap T)+f_{\text{prob}}(S\cup T)$ .
@@
$\tilde{C}^{*,S\cup T}_{m}={C}^{*,S\cup T}_{m}\setminus\{i\in{C}^{*,S\cup T}_{m}:\text{play $i$ doesn't get resource}\},$
@@
[…]
@@
The first inequality follows by Lemma (1); the second inequality follows by $1-\alpha(|S^{*}|)\leq 1$; the third inequality follows by and Lemma 2; the fourth inequality follows by [24], which states that a simple greedy algorithm with an objective function that is monotone and submodular can obtain an approximation factor of an approximation factor of 11/e1-1/e. Since $f_{\text{prob}}(S)$ is monotonely increasing and submodular shown in Lemma 3 and Lemma 4 respectively, we have the approximation factor 11/e1-1/e. The next one is the second equality, and it follows by $|S_{i}|=i$ for all $0\leq i\leq I$; the fifth inequality by the fact that $\tilde{S}^{*}$ maximizes $(1-\alpha(|S_{i}|))f\left(S_{i}\right)$; the last inequality follows by Lemma 5. ∎
@@ Appendix 3 Proofs for the online case
[…]
@@
[…]
@@
where $\delta\in(0,1)$, $\epsilon^{(t)}_{m,k}=\sqrt{\left(1+{n^{(t)}_{m,k}}\right)\frac{\ln\left(\sqrt{n^{(t)}_{m,k}+1}/\delta\right)}{2\left(n^{(t)}_{m,k}\right)^{2}}}$ if $n^{(t)}_{m,k}>0$ and $\epsilon^{(t)}_{m,k}$$\infty$ if $n^{(t)}_{m,k}=0$.
@@
$-\mathbb{E}\left[\max_{{\mathbf{C}}_{t}}\mathcal{R}^{\text{total}}(S_{t},\{N_{t,m},{\mathbf{X}}_{t,m}\}_{m\in S_{t}},\{{\mathbf{p}}_{t,m},{\bm{\mu}}_{m}\}_{m\notin S_{t}},{\mathbf{C}}_{t})\right],\,\text{(by Definition)},$
@@
  • Environment $\mathcal{E}_{0}$: All $M$ arms follow an identical Bernoulli($\mu$) distribution.
  • Environment $\mathcal{E}_{1}$: Arm 1 is Bernoulli($\mu$), while the remaining M1M-1 arms are Bernoulli($\mu-\Delta$), where $\Delta=c/\sqrt{T}$ for some constant $c>0$.
@@
Using results from Le Cam’s method and Pinsker’s inequality [5, 17], the total variation distance between Bernoulli($\mu$) and Bernoulli($\mu-\Delta$) is bounded by $\Delta$. To reliably distinguish between $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$, the algorithm needs at least $\Omega(1/\Delta^{2})$ samples per arm. Setting $\Delta=c/\sqrt{T}$ leads to $\Omega(T)$ samples across all arms, which is infeasible within $T$ rounds.